40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun May 16 12:51:08 2010, kiwamu, Update, Green Locking, reflected beam at PD 
    Reply  Sun May 16 19:25:45 2010, Koji, Update, Green Locking, reflected beam at PD 
       Reply  Thu May 20 13:29:40 2010, kiwamu, Update, Green Locking, mode profile at PPKTP crystal PPKTPmode.png
          Reply  Thu May 20 14:18:59 2010, kiwamu, Update, Green Locking, mode profile of 40m cavity MMT40mcavity.png
             Reply  Wed May 26 04:14:21 2010, kiwamu, Update, Green Locking, locked OLTF_endPDH.png
                Reply  Wed May 26 14:38:02 2010, Koji, Update, Green Locking, locked 
                   Reply  Thu May 27 02:22:24 2010, kiwamu, Update, Green Locking, more details stimes.pngspectrum.pngETMX_small.png
                      Reply  Thu May 27 09:43:50 2010, rana, Update, Green Locking, more details 
                         Reply  Mon Jul 12 13:16:53 2010, kiwamu, Update, Green Locking, PZT feedback at X end arm_day_night.png
                      Reply  Wed Jun 9 11:13:11 2010, kiwamu, Update, Green Locking, lock with PDH box 
                         Reply  Fri Jun 11 11:54:42 2010, kiwamu, Update, Green Locking, end PDH with thermal feedback LPF.pngthermal_feedback.png
                            Reply  Fri Jun 11 13:32:28 2010, Koji, Update, Green Locking, end PDH with thermal feedback 
                         Reply  Sun Jun 20 18:28:25 2010, kiwamu, Update, Green Locking, Re: lock with PDH box 
                Reply  Wed Jul 21 19:11:18 2010, Dmass, Update, Green Locking, locked 
Message ID: 3195     Entry time: Mon Jul 12 13:16:53 2010     In reply to: 2999
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: PZT feedback at X end 

The feedback signal going to the laser PZT at the X end station was measured in the daytime and the nighttime.

It's been measured while the laser frequency was locked to the arm cavity with the green light.

arm_day_night.png

 

The red curve was measured at 3pm of 8/July Friday. And the blue curve was measured at 12am of 9/July Saturday. 

As we can see on the plot, the peak-peak values are followers

              daytime:  ~ 4Vpp

       nighttime:  ~1.8Vpp

It is obvious that the arm cavity is louder in the daytime by a factor of about 2.

Note: the feedback signal is sent to the PZT only above 1Hz because the low frequency part is stabilized mostly by the crystal temperature (see this entry)

Quote:

 What we care about is the peak-peak value of the PZT feedback signal measured on a scope for ~30 seconds.

ELOG V3.1.3-